Fermi National Laboratory


Accelerator Update - July 16- July 18

The integrated luminosity for last week was 8467.4 nb-1.

Wednesday July 16
The day shift began with Operations putting the Tevatron (TeV) into shot setup, and with MiniBooNE taking beam.

Operations reported having trouble with the TeV reverse injection at 9:34 AM. Experts came in to investigate. Operations put shot setup on hold while this and other problems were checked out. Around 3 PM, power supply technicians discovered and replaced a bad current transformer on T:SD.

All the Main Injector (MI)RF stations went down at 2:32 PM. Operations called experts in to investigate.

The Antiproton Source (Pbar) resumed stacking at 2:51 PM.

Operations put the TeV through a dry squeeze at 2:58 PM and then started shot setup at 3:15 PM.

Operations tuned the Booster accelerator at 4:30 PM.

At 5:28 PM, Operations established store 2792 with an initial luminosity of 30.6E30.

Operators noticed, at 6:19 PM, that MI LCW pressures had been dropping. Experts thought that a nitrogen gas bottle might be empty and that the gas bottle's regulator might be faulty. Operators changed the gas bottle and techs came in to change the regulator.

At 6:30 PM, the TeV suffered a sector E2 quench. The store was lost. After an investigation, the cause was thought to be due to a small vacuum leak.

The temperatures for the cryo system at E2 were still rising at 7:28 PM. Operations contacted a cryo expert for help.

Operators made the TeV ready for access at 7:48 PM and changed some equipment in sector E2. They were done by 10:28 PM. The vacuum techs arrived and had two vacuum pumps installed by 11:31 PM.

Operations started a TeV checkout at 11:41 PM.

Thursday July 17
The midnight shift began with the TeV still recovering from the quench, with Pbar stacking, and with MiniBooNE taking beam.

Operations put the TeV into shot setup at 3:28 AM.

The Pbar sequencer began causing problems 4:46 AM. Experts came in. By 7:03 AM, they had found and replaced a bad MADC power supply.

At 4:51 AM, the TeV suffered a quench at sector F1. This occurred while operators were trying to set up the Pbarhelix for a chromaticity measurement.

Operations put the TeV into shot setup and at 8:55 AM established store 2795 with an initial luminosity of 22.88E30. Pbar resumed stacking.

At 5:06 PM, a rolling thunderstorm paused over the Fermi accelerators just long enough to create a lightning induced power glitch. This caused store 2795 to abort and the stack to disappear. Although many things tripped off all over the complex, the Recycler managed to maintain its store of protons.

Operations called in TeV RF techs to help fix three RF stations that went down hard. They had the stations back up by 5:30 PM.

Operations reestablished beam to MiniBooNE at 5:32 PM. Stacking resumed at 6 PM.

A MiniBooNE expert conducted a target scan at 9:11 PM.

Friday July 18
The midnight shift began with Pbar stacking, and with TeV and MiniBooNE experts conducting studies.

MiniBooNE experts completed their studies at 12:52 AM, and resumed normal running.

A TeV RF station's (TRF1) water heater tripped off at 3:23 AM. TeV studies continued. Operators called in an expert at 4:45 AM.

Operations put the TeV into a dry squeeze at 7:43 AM.

Today's Plan
The plan for today and this weekend is to stack and store. There may be some studies conducted in between stores.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 7/18/2003   email Fermilab